# Uncertainty Modeling in Graph Neural Networks via Stochastic Differential Equations

### v4

@@ Abstract
We propose a novel Stochastic Differential Equation (SDE) framework to address the problem of learning uncertainty-aware representations for graph-structured data. While Graph Neural Ordinary Differential Equations (GNODEs) have shown promise in learning node representations, they lack the ability to quantify uncertainty. To address this, we introduce Latent Graph Neural Stochastic Differential Equations (LGNSDE), which enhance GNODE by embedding randomness through a Bayesian prior-posterior mechanism for epistemic uncertainty and Brownian motion for aleatoric uncertainty. By leveraging the existence and uniqueness of solutions to graph-based SDEs, we prove that the variance of the latent space bounds the variance of model outputs, thereby providing theoretically sensible guarantees for the uncertainty estimates. Furthermore, we show mathematically that LGNSDEs are robust to small perturbations in the input, maintaining stability over time. Empirical results across several benchmarks demonstrate that our framework is competitive in out-of-distribution detection, robustness to noise, and active learning, underscoring the ability of LGNSDEs to quantify uncertainty reliably.
@@ 1 Introduction
To address this gap, we propose Latent Graph Neural Stochastic Differential Equations (LGNSDE), a method that perturbs features during both the training and testing phases using Brownian motion noise, allowing for handling noise and aleatoric uncertainty. We assume a prior SDE on the latent space and learn a posterior SDE using a GNN as the drift function. This Bayesian approach to the latent space allows us to quantify epistemic uncertainty. As a result, our model can capture and quantify both epistemic and aleatoric uncertainties. More specifically, our contributions are as follows:
@@ 3 Methodology
Inspired by Graph Neural ODEs (Poli et al., 2019) and Latent SDEs (Li et al., 2020), we now introduce our model: Latent Graph Neural SDEs -- LGNSDEs (Figure 1), which use SDEs to define prior and approximate posterior stochastic trajectories for $\mathbf{H}(t)$ (Xu et al., 2022). Furthermore, LGNSDEs can be viewed as the continuous representations of existing discrete architectures (A.4).
@@ 4 Theoretical Guarantees
[…]
@@
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ is a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the LGNSDE with these initial conditions remains bounded across time $t$ [^2] Note that while the bound is exponential in $t$, in practice, the time horizon is usually constrained to a limited range, such as $t\in[0,1]$. Within this interval, the exponential factor does not grow excessively, ensuring that the deviation between the perturbed and unperturbed solutions remains under control., specifically
@@ 5.2 Out of Distribution Detection
The bottom of Figure 2 shows the test entropy distribution for in-distribution (blue) and out-of-distribution (red) data. For each test sample, predictions were made over $C-1$ classes, excluding the left-out class. The OOD class exhibits higher entropy, indicating greater uncertainty. While most models show similar entropy distributions for both data types, our LGNSDE model achieves a clear separation, with a 50% higher mean entropy for OOD data compared to in-distribution data. Other models show less than a 10% difference between the two distributions.
@@ A.3 Proofs
[…]
@@
[…]
@@
[…]
@@ C.3 Further Benchmarking and Complexities
$N$ Number of nodes in the graph

### v5

@@ Abstract
We propose a novel Stochastic Differential Equation (SDE) framework to address the problem of learning uncertainty-aware representations for graph-structured data. While Graph Neural Ordinary Differential Equations (GNODEs) have shown promise in learning node representations, they lack the ability to quantify uncertainty. To address this, we introduce Latent Graph Neural Stochastic Differential Equations (LGNSDE), which enhance GNODE by embedding randomness through a Bayesian prior-posterior mechanism for epistemic uncertainty and Brownian motion for aleatoric uncertainty. By leveraging the existence and uniqueness of solutions to graph-based SDEs, we prove that the variance of the latent space bounds the variance of model outputs, thereby providing theoretically sensible guarantees for the uncertainty estimates. Furthermore, we show mathematically that LGNSDEs are robust to small perturbations in the input, maintaining stability over time. Empirical results across several benchmarks demonstrate that our framework is competitive in out-of-distribution detection, robustness to noise, and active learning, underscoring the ability of LGNSDEs to quantify uncertainty reliably. Code is available at github.com/Richard-Bergna/GraphNeuralSDE.
@@ 1 Introduction
To address this gap, we propose Latent Graph Neural Stochastic Differential Equations (LGNSDE) [^1] Code available at https://github.com/Richard-Bergna/GraphNeuralSDE, a method that perturbs features during both the training and testing phases using Brownian motion noise, allowing for handling noise and aleatoric uncertainty. We assume a prior SDE on the latent space and learn a posterior SDE using a GNN as the drift function. This Bayesian approach to the latent space allows us to quantify epistemic uncertainty. As a result, our model can capture and quantify both epistemic and aleatoric uncertainties. More specifically, our contributions are as follows:
@@ 3 Methodology
Inspired by Graph Neural ODEs (Poli et al., 2019) and Latent SDEs (Li et al., 2020), we now introduce our model: Latent Graph Neural SDEs - LGNSDEs (Figure 1), which use SDEs to define prior and approximate posterior stochastic trajectories for $\mathbf{H}(t)$ (Xu et al., 2022). Furthermore, LGNSDEs can be viewed as the continuous representations of existing discrete architectures (A.4).
@@ 4 Theoretical Guarantees
$\mathbf{H}_{i}(t)=e^{t\Psi}\mathbf{H}_{i}(0)+\int_{0}^{t}e^{(t-s)\Psi}\mathbf{G}_{\mathcal{G}}(\mathbf{H}_{i}(s))d\mathbf{W}(s),$
@@
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ is a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the LGNSDE with these initial conditions remains bounded across time tt [^3] Note that while the bound is exponential in $t$, in practice, the time horizon is usually constrained to a limited range, such as $t\in[0,1]$. Within this interval, the exponential factor does not grow excessively, ensuring that the deviation between the perturbed and unperturbed solutions remains under control., specifically
@@ 5.2 Out of Distribution Detection
The bottom of Figure 2 shows the test entropy distribution for in-distribution (blue) and out-of-distribution (red) data. For each test sample, predictions were made over C1C-1 classes, excluding the left-out class. The OOD class exhibits higher entropy, indicating greater uncertainty. While most models show similar entropy distributions for both data types, our LGNSDE model achieves a clear separation, with a 50% higher mean entropy for OOD data compared to in-distribution data. Other models show less than a 10% difference between the two distributions.
@@ A.3 Proofs
$\mathbb{E}\left[\left\|\int_{0}^{t}\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)d\mathbf{W}(u)\right\|_{F}^{2}\right]=\mathbb{E}\left[\int_{0}^{t}\|\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)\|_{F}^{2}du\right].$
@@
$\text{Var}(\mathbf{\hat{y}}(t))=\int_{0}^{t}\text{tr}\left(\mathbf{h}^{\prime}(\mathbf{H}(u))^{\top}\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)^{\top}\mathbf{h}^{\prime}(\mathbf{H}(u))\right)du.$
@@
$\text{Var}(\mathbf{y}(t))\leq L_{h}^{2}\int_{0}^{t}\|\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)\|_{F}^{2}du=L^{2}_{h}\text{Var}(\mathbf{H}(t)).$
@@ C.3 Further Benchmarking and Complexities
NN Number of nodes in the graph